Economics question.
Developing the cash flow for each alternative in a study is a pivotal, and usually the most difficult, step in the engineering economic analysis procedure. An integrated approach for developing cash flows includes three major components: (1) a work breakdown structure (WBS) definition of the project, (2) a cost and revenue structure that identifies all the cost and revenue elements involved in the study, and (3) estimating techniques (models). Chapter 4 presented the fundamental time value of money relationships. Discuss the concept of equivalence - if two cash flows (or a series of cash flows) are equivalent for a stated interest rate, and under what circumstances you would be willing to trade one for the other. Search entries or author Unread Reply Replies are only visible to those who have posted at least one reply * Previous Next ch
Answer #1

If two cash flows are equivalent, I would be willing to select one over the other if the cash flows of this particular project tare occurring earlier than the other one. For instance if Project A yields cash flows of $100 per annum for 5 years starting next year, its present value at 10% rate will be 100*P/A,10%,5 = 100*3.791 = $379.1

On the other hand if Project A yields the same cash flows of$100 for 5 years but starting after 2 years, its present value at 10% will be 100*P/A,10%,5 /(1+r)^1 = 100*3.791/1.1^1 =$ 344.64

Hence we would be better off selecting Project A since it has higher present worth.